You own Nails Solutions, specializing in premium manicures.

You want to analyze your direct labor costs.
Based on the standards, each and every manicure should take 1/2 hour (30 minutes)
Last week your employees did 200 manicures.
You pay your employees by the hour and last week you paid them a total of $1,840 for 140 hours
of work.

The standard rate of pay is $12 per hour.

1. what is the labor rate variance? Show formula

2. What was the weeks labor efficiency variance? Show formula
Answer #1

Standard Rate = $ 12 per hour

Standard hours (200 manicures * 30 minutes ) = 100 hours

Actual Rate ($ 1840/140) = $ 13.14 hour

Actual hour = 140 hours

1. Labour rate variance = Actual hours ( Standard Rate - Actual Rate )

= 140 (12 - 13.14) = 160 (A)

2. Labour efficiency variance = Standard rate (standard hours - actual hours )

= 12 (100 -140) = 480 (A)